# Highest-weight vectors and three-point functions in GKO coset decomposition

### v1

@@ Abstract
We revisit the classical Goddard-Kent-Olive coset construction. We find the formulas for the highest weight vectors in coset decomposition and calculate their norms. We also derive formulas for matrix elements of natural vertex operators between these vectors. This leads to relations on conformal blocks. Due to the AGT relation, these relations are equivalent to blowup relations on Nekrasov partition functions with the presence of the surface defect. These relations can be used to prove Kyiv formulas for the Painlevé tau-functions (following Nekrasov’s method).
@@ 1 Introduction
Goddard Kent Olive coset construction is one of the most basic and important constructions in the theory of vertex algebras. This construction can be viewed as an affine analog of the decomposition of tensor product of representations of $\mathfrak{sl}(2)$. Let $\widehat{\mathfrak{sl}}(2)_{k}$ denotes affine (vertex) algebra $\mathfrak{sl}_{2}$ where central element acts by number $k\in\mathbb{C}$. On the tensor product of representations of $\widehat{\mathfrak{sl}}(2)_{1}$ and $\widehat{\mathfrak{sl}}(2)_{k+1}$ there is a natural action of diagonal $\widehat{\mathfrak{sl}}(2)_{k+1}$. It was observed in [GKO86] that multiplicity spaces have a natural structure of representation of the Virasoro algebra with central charge $c=13-6(\frac{k+2}{k+3}+\frac{k+3}{k+2})$. Moreover, there is an isomorphism of vertex algebras (assume that $k\not\in\mathbb{Q}$)
@@
One of the approaches to DOZZ formula [Tes95] is based on the interplay of two Virasoro algebras with the same central charge in Liouville theory, namely chiral and antichiral algebras. Similarly in the proof of Theorem 4.5 we used interplay of $\widehat{\mathfrak{sl}}(2)_{k+1}$ and $\mathrm{Vir}_{c}$ symmetries, and relation between $k$ and $c$ is very important. It appears that right side of (1.1) is a simplest case of corner vertex algebras $\mathcal{D}_{n,k}^{\psi}$ [CG20], [CDGG21]. We hope that methods used in this paper can be used to study more general corner algebras, at least corresponding to $\mathfrak{sl}(2)$. Another possible generalization is the study of generic $\mathfrak{sl}(2)$ coset of the form $\mathfrak{sl}(2)_{k_{1}}\oplus\mathfrak{sl}(2)_{k_{1}}/\mathfrak{sl}(2)_{k_{1}+k_{2}}$. In the last case, the coset algebra is $\mathcal{D}(2|1,\alpha)$ $W$-algebra [FS01].
@@
Due to AGT correspondence the relation (1.2) has interpretation in terms of 4d supersymmetric gauge theory. Namely, the function $\mathrm{F}_{c}$ is equal (again, up to elementary function) to Nekrasov partition function for $SU(2)$ gauge theory and function $\Psi_{k}$ is equal to Nekrasov partition function with surface defect. The relation (1.2) itself takes the meaning of the blowup relation [NY05], to be more precise these are blowup relations with the presence of surface defect suggested in [Nek24], [JN20]. Hence our Theorem 4.5 gives the proof of blowup relations with surface defect (for $SU(2)$ group). Such method of the proof of blowup relations follows [BFL16], see also [ACF22].
It is worth mentioning that coset (or blowup) relations (1.2) have remarkable application in the theory of isomonodromic deformations as was shown in [Nek24], [JN20]. In CFT language this idea can be restated as a $k\rightarrow\infty$ limit of algebras (1.1) or conformal blocks (1.2). The algebra $\widehat{\mathfrak{sl}}(2)_{k+1}$ in this limit becomes classical, so we get vertex algebra with big center [Fei17], [CDGG21], [FL22]. The spectrum of the center is the space of $\mathfrak{sl}(2)$ connections. The conformal blocks $\Psi_{k}$ satisfy Knizhnik-Zamolodchikov equations which in classical limit goes to isomonodromic deformation equations [Res92], [Har96], [Nek24]. Hence one got a proof of Kyiv formula for isomonodromic tau function [GIL12].
@@
In section 3 we revisit GKO coset construction. The first result is an explicit formula for the highest weight vectors $u_{l}(\lambda)$, see Theorem 3.7. This formula is written in free field realization, i.e. we consider Wakimoto representation of $\widehat{\mathfrak{sl}}(2)_{k}$.  [^1] While this paper was in preparation, the authors became aware of the paper [HR23] which has some overlap with our results. In particular, a formula for the vectors $u_{l}(\lambda)$ in free field realization was found there, which is different from our formula. It would be interesting to compare these formulas. The existence of the formula in free field realization is a standard feature in the representation theory of vertex algebras, c.f. formulas for Virasoro singular vectors in terms of Jack symmetric functions [MY95] or formulas for the highest weight vector in decomposition [BBLT13].
@@
Section 4 is devoted to the proof of the Theorem 4.5. The idea is to consider a four-point conformal block with an insertion of a degenerate field $b(x,z)$. Essentially, we mimic a standard approach to study theories with chiral and antichiral symmetries [ZZ89], [Tes95], [Tes99], in purely chiral setting this also used in [BS15]. The inserted field $b(x,z)$ in terms of left side of (1.1) is a product of operator of spin $\frac{1}{2}$ for $\widehat{\mathfrak{sl}}(2)_{1}$ and identity for $\widehat{\mathfrak{sl}}(2)_{k}$, while in terms of the right side of (1.1) is a product of spin $\frac{1}{2}$ for $\widehat{\mathfrak{sl}}(2)_{k+1}$ and $\Phi_{1,2}$ for $\mathrm{Vir}_{c}$.
@@ Acknowledgements
M.B. is grateful to Nikita Nekrasov for explanations about relation between blowup relations with surface defect and Kyiv formula in 2013, which eventually lead to this work. We are grateful to T. Creutzig, N. Genra, M. Lashkevich, A. Litvinov, H. Nakajima, N. Nekrasov, M. Noumi, F. Petrov, A. Shchechkin, J. Teschner, A. Zamolodchikov for useful discussions. The results of the paper were reported at conferences and seminars in Tokyo (October 2022, January 2023), Edinburgh (September 2023), Trieste (December 2023), Moscow/Zoom (December 2023), Hamburg/Toronto/Zoom (March 2023); we are grateful to the organizers and participants for their interest and remarks.
The work of M.B is partially supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme under grant agreement No 948885. M.B. is very grateful to Kavli IPMU and especially Y. Fukuda, M. Kapranov, K. Kurabayashi, H. Nakajima, A. Okounkov, T. Shiga, K. Vovk, for the hospitality during 2022-2023 years.
@@ 2.1 Definition
Let $\widehat{\mathfrak{sl}}(2)=\mathfrak{sl}(2)\otimes\mathbb{C}[[t,t^{-1}]\oplus\mathbb{C}K$ denotes the central extension of the of the algebra of Lawrent power series with coefficients in $\mathfrak{sl}_{2}$. It has topological basis $e_{m}=e\otimes t^{m}$, $h_{m}=h\otimes t^{m}$, $f_{m}=f\otimes t^{m}$, $K$ with commutation relations
@@
Verma module $\mathcal{M}_{\lambda,k}$ is a module over algebra $\widehat{\mathfrak{sl}}(2)$ which is generated freely by $e_{n},h_{n},f_{n+1}$, $(n\in\mathbb{Z}_{<0})$ acting on a vector highest weight vector $v_{\lambda,k}$ such that
@@
Let $\widehat{\mathfrak{b}}=\langle K,e_{n},h_{n},f_{n+1}|n\in\mathbb{Z}_{\geq 0}\rangle$ denotes Borel subalgebra in $\widehat{\mathfrak{sl}}(2)$. The formulas (2.3), (2.4) define one dimensional representation of $\widehat{\mathfrak{b}}$, which we denote by $\mathbb{C}_{\lambda,k}$. The Verma module can be also written as an induced module $\mathcal{M}_{\lambda,k}=\operatorname{Ind}^{\widehat{\mathfrak{sl}}(2)}_{\widehat{\mathfrak{b}}}\mathbb{C}_{\lambda,k}$.
@@
We will call pair $(\lambda,k)$ generic if they are linearly independent over $\mathbb{Q}$. It is typically sufficient to require that they do not satisfy conditions (2.5). Similarly, we call $k$ generic if it is irrational.
@@ 2.3 Sugawara construction
We will use normal ordering of two fields. See, e.g. [FBZ04, Ch. 2]
@@ 2.4 Integrable modules on level $1$
The Fock module $\mathcal{F}_{\alpha}$ is a Heisenberg algebra module freeley generated by $a_{n}$, $n<0$ acting on a highest weight vector $|\alpha\rangle=v_{\frac{\alpha}{\sqrt{2}}}$ such that
@@
The modules $\mathcal{L}_{0,1}$, $\mathcal{L}_{1,1}$ are integrable [Kac90]. The highest weight vectors $v_{n}\in\mathcal{F}_{\sqrt{2}n}$, $n\in\frac{1}{2}\mathbb{Z}$ in components from decompositions (2.21) are called extremal vectors. It follows from the Theorem 2.15 that $v_{0},v_{1/2}$ are highest weight vectors of $\mathcal{L}_{0,1}$ and $\mathcal{L}_{1,1}$ correspondingly and other extremal vectors can be found by the formulas
@@
The following formulas for characters follows from the Theorem 2.15.
@@ 2.5 Wakimoto module
The $\widehat{\mathfrak{sl}}(2)_{k}$-module $\mathcal{F}_{\lambda,k}$ is called Wakimoto module. See also [FBZ04, Ch.11-12]. The following proposition easily follows from the equality of the characters.
@@ 2.6 Vertex operators
The are (at least) two standard approaches to the definition of vertex operators in conformal field theory. The first one is based on the definition of conformal blocks via coinvariants and operator-state correspondence. In the second approach vertex operators are defined using commutation relation with the algebra generators. We will use both approaches. We start from the first one, mainly following the paper [FM94]. We restrict ourselves to the genus zero conformal block. We fix a global coordinate $t$ on $\mathbb{P}^{1}$.
@@ 2.6.1 Generic representations
This subalgebra is conjugated from $\mathfrak{b}=\langle h,e\rangle$ by the $e^{xf}\in SL(2)$. Let $h_{x},e_{x}$ be generators of $\mathfrak{b}_{x}$ such that $[h_{x},e_{x}]=2e_{x}$, $e_{x}\in[\mathfrak{b}_{x},\mathfrak{b}_{x}]$. For example for one can take $h_{x}=e^{xf}he^{-xf}$, $e_{x}=e^{xf}ee^{-xf}$ for $x\neq\infty$ and $h_{\infty}=-h$, $e_{\infty}=-f$ for $x=\infty$.
@@
$a_{x,z}:\widehat{\mathfrak{sl}}(2)\rightarrow\widehat{\mathfrak{sl}}(2)_{z},$ (2.31)
@@
Of course these formulas do not work if $x$ or $z$ are equalt to infinity. So in the neighbourhood of (,)($\infty$,$\infty$)( ∞ , ∞ ) we use another isomorphism
@@
We use the same notation for the corresponding maps between Verma modules $a_{x,z}:\mathcal{M}_{\nu,k}\rightarrow\mathcal{M}_{\nu,k}(x,z)$ mapping $v_{\nu,k}\mapsto v_{\nu,k}(x,z)$.
@@
Let $x_{1},x_{2},x_{3}\in\mathbb{P}^{1}$ and $x_{1},x_{2},x_{3}\in\mathbb{P}^{1}$ be tuples of different points. There is a unique up to constant homomorphism of $\mathfrak{sl}(2)\otimes\mathbb{C}[\mathbb{P}^{1}\setminus\{\vec{z}\}]$-modules
@@
We define map $\mathtt{m}_{k}(\vec{x},\vec{z})$ as a composition $\mathtt{m}_{k}(\vec{x},\vec{z})=\mathtt{m}_{k}\circ(a_{x_{1},z_{1}}\otimes a_{x_{2},z_{2}}\otimes a_{x_{3},z_{3}})$
@@
Using $PGL(2)$ action on $\mathbb{P}^{1}$ (global conformal transformations) we can assume that $\vec{z}=\{0,z,\infty\}$, and similarly we can assume that $\vec{x}=\{0,x,\infty\}$ using global $SL(2)$ acting on $\mathbb{C}[\mathbb{P}^{1}\setminus\vec{z}]$. We write simply
@@
The following Proposition follows from a direct computation (cf. [FM94, Prop. 3.1]). On the other hand this proposition can be considered as a definition of the vertex operator, (see [AY92]).
In order to write down $\mathtt{Y}(v|x,z)$ for arbitrary $v\in\mathcal{M}_{\nu,k}$ we will need the notations
$e(x,z)=e^{xf_{0}}e(z)e^{-xf_{0}}=e(z)-xh(z)-x^{2}f(z),$ (2.40)
$h(x,z)=e^{xf_{0}}h(z)e^{-xf_{0}}=h(z)+2xf(z),~{}~{}f(x,z)=e^{xf_{0}}f(z)e^{-xf_{0}}=f(z).$ (2.41)

### v3

@@ Abstract
We revisit the classical Goddard-Kent-Olive coset construction. We find the formulas for the highest weight vectors in coset decomposition and calculate their norms. We also derive formulas for matrix elements of natural vertex operators between these vectors. This leads to relations on conformal blocks. Due to the AGT correspondence, these relations are equivalent to blowup relations on Nekrasov partition functions with the presence of the surface defect. These relations can be used to prove Kyiv formulas for the Painlevé tau-functions (following Nekrasov’s method).
@@ 1 Introduction
Goddard-Kent-Olive coset construction is one of the most basic and important constructions in the theory of vertex algebras. This construction can be viewed as an affine analog of the decomposition of tensor product of representations of $\mathfrak{sl}(2)$. Let $\widehat{\mathfrak{sl}}(2)_{k}$ denotes affine (vertex) algebra $\mathfrak{sl}_{2}$ where central element acts by number $k\in\mathbb{C}$. On the tensor product of representations of $\widehat{\mathfrak{sl}}(2)_{1}$ and $\widehat{\mathfrak{sl}}(2)_{k+1}$ there is a natural action of diagonal $\widehat{\mathfrak{sl}}(2)_{k+1}$. It was observed in [GKO86] that multiplicity spaces have a natural structure of representation of the Virasoro algebra with central charge $c=13-6(\frac{k+2}{k+3}+\frac{k+3}{k+2})$. Moreover, there is an isomorphism of vertex algebras (assume that $k\not\in\mathbb{Q}$)
@@
One of the approaches to DOZZ formula [Tes95] is based on the interplay of two Virasoro algebras with the same central charge in Liouville theory, namely chiral and antichiral algebras. Similarly, in the proof of Theorem 4.5 we use interplay of $\widehat{\mathfrak{sl}}(2)_{k+1}$ and $\mathrm{Vir}_{c}$ symmetries in which relation between $k$ and $c$ is very important. It appears that right side of (1.1) is a simplest case of corner vertex algebras $\mathcal{D}_{n,k}^{\psi}$ [CG20], [CDGG21]. We hope that methods used in this paper can be used to study more general corner algebras, at least corresponding to $\mathfrak{sl}(2)$. Another possible generalization is the study of generic $\mathfrak{sl}(2)$ coset of the form $\mathfrak{sl}(2)_{k_{1}}\oplus\mathfrak{sl}(2)_{k_{1}}/\mathfrak{sl}(2)_{k_{1}+k_{2}}$. In the last case, the coset algebra is $\mathcal{D}(2|1,\alpha)$ $W$-algebra [FS01].
@@
Due to AGT correspondence the relation (1.2) has interpretation in terms of 4d supersymmetric gauge theory. Namely, the function $\mathrm{F}_{c}$ is equal (again, up to elementary function) to Nekrasov partition function for $SU(2)$ gauge theory and function $\Psi_{k}$ is equal to Nekrasov partition function with surface defect. The relation (1.2) itself takes the meaning of the blowup relation [NY05], to be more precise this is the blowup relation with the presence of surface defect suggested in [Nek24], [JN20]. Hence our Theorem 4.5 gives the proof of blowup relations with surface defect (for $SU(2)$ group). Such method of the proof of blowup relations follows [BFL16], see also [ACF22].
It is worth mentioning that coset (or blowup) relations (1.2) have remarkable application in the theory of isomonodromic deformations as was shown in [Nek24], [JN20]. In CFT language this idea can be restated as a $k\rightarrow\infty$ limit of algebras (1.1) or conformal blocks (1.2). The algebra $\widehat{\mathfrak{sl}}(2)_{k+1}$ in this limit becomes classical, so we get vertex algebra with big center [Fei17], [CDGG21], [FL24]. The spectrum of the center is the space of $\mathfrak{sl}(2)$ connections. The conformal blocks $\Psi_{k}$ satisfy Knizhnik-Zamolodchikov equations which in classical limit goes to isomonodromic deformation equations [Res92], [Har96], [Nek24]. Hence one got a proof of Kyiv formula for isomonodromic tau function [GIL12].
@@
In section 3 we revisit GKO coset construction. The first result is an explicit formula for the highest weight vectors $u_{l}(\lambda)$, see Theorem 3.7. This formula is written in free field realization, i.e. we consider Wakimoto representation of $\widehat{\mathfrak{sl}}(2)_{k}$.  [^1] While this paper was in preparation, the authors became aware of the paper [HR25] which has some overlap with our results. In particular, a formula for the vectors $u_{l}(\lambda)$ in free field realization was found there, which is different from our formula. It would be interesting to compare these formulas. The existence of the formula in free field realization is a standard feature in the representation theory of vertex algebras, c.f. formulas for Virasoro singular vectors in terms of Jack symmetric functions [MY95] or formulas for the highest weight vector in decomposition [BBLT13].
@@
Section 4 is devoted to the proof of the Theorem 4.5. The idea is to consider a four-point conformal block with an insertion of a degenerate field $b(x,z)$. Essentially, we mimic a standard approach to study theories with chiral and antichiral symmetries [ZZ89], [Tes95], [Tes99], in purely chiral setting this was also used in [BS15]. The inserted field $b(x,z)$ in terms of left side of (1.1) is a product of operator of spin $\frac{1}{2}$ for $\widehat{\mathfrak{sl}}(2)_{1}$ and identity for $\widehat{\mathfrak{sl}}(2)_{k}$, while in terms of the right side of (1.1) is a product of spin $\frac{1}{2}$ for $\widehat{\mathfrak{sl}}(2)_{k+1}$ and $\Phi_{1,2}$ for $\mathrm{Vir}_{c}$.
@@ Acknowledgements
M.B. is grateful to Nikita Nekrasov for explanations about relation between blowup relations with surface defect and Kyiv formula in 2013, which eventually lead to this work. We are grateful to T. Creutzig, N. Genra, M. Lashkevich, A. Litvinov, H. Nakajima, N. Nekrasov, M. Noumi, F. Petrov, A. Shchechkin, J. Teschner, A. Zamolodchikov for useful discussions. The results of the paper were reported at conferences and seminars in Tokyo (October 2022, January 2023), Edinburgh (September 2023), Trieste (December 2023), Moscow/Zoom (December 2023), Hamburg/Toronto/Zoom (March 2024); we are grateful to the organizers and participants for their interest and remarks.
The work of M.B is partially supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme under grant agreement No 948885. M.B. is very grateful to Kavli IPMU and especially Y. Fukuda, M. Kapranov, K. Kurabayashi, H. Nakajima, A. Okounkov, T. Shiga, K. Vovk, for the hospitality during 2022–2023 years.
@@ 2.1 Definition
Let $\widehat{\mathfrak{sl}}(2)=\mathfrak{sl}(2)\otimes\mathbb{C}[[t,t^{-1}]\oplus\mathbb{C}K$ denotes the central extension of the of the algebra of Laurent series with coefficients in $\mathfrak{sl}_{2}$. It has topological basis $e_{m}=e\otimes t^{m}$, $h_{m}=h\otimes t^{m}$, $f_{m}=f\otimes t^{m}$ $(m\in\mathbb{Z})$, and $K$ with commutation relations
@@
Verma module $\mathcal{M}_{\lambda,k}$ is a module over algebra $\widehat{\mathfrak{sl}}(2)$ which is generated freely by $e_{n},h_{n},f_{n+1}$, $(n\in\mathbb{Z}_{<0})$ acting on a highest weight vector $v_{\lambda,k}$ such that
@@
Let $\widehat{\mathfrak{b}}=\langle K,e_{n},h_{n},f_{n+1}|n\in\mathbb{Z}_{\geq 0}\rangle$ denotes Borel subalgebra in $\widehat{\mathfrak{sl}}(2)$ (note that $K,e_{n},h_{n},f_{n+1}$ form a topological generating set of $\widehat{\mathfrak{b}}$, i.e. infinite sums are allowed). The formulas (2.3), (2.4) define one dimensional representation of $\widehat{\mathfrak{b}}$, which we denote by $\mathbb{C}_{\lambda,k}$. The Verma module can be also written as an induced module $\mathcal{M}_{\lambda,k}=\operatorname{Ind}^{\widehat{\mathfrak{sl}}(2)}_{\widehat{\mathfrak{b}}}\mathbb{C}_{\lambda,k}$.
@@
We will call pair $(\lambda,k)$ generic if $\lambda$ and $k$ are linearly independent over $\mathbb{Q}$. It is typically sufficient to require that they do not satisfy conditions (2.5). Similarly, we call $k$ generic if it is irrational.
@@ 2.3 Sugawara construction
We will use normal ordering of fields. See, e.g. [FBZ04, Ch. 2]
@@ 2.4 Integrable modules on level $1$
The Fock module $\mathcal{F}_{\alpha}$ is a Heisenberg algebra module freely generated by $a_{n}$, $n<0$ acting on a highest weight vector $|\alpha\rangle=v_{\frac{\alpha}{\sqrt{2}}}$ such that
@@
Here and below in exponents of Heisenberg algebra we use bosonic normal ordering.
The modules $\mathcal{L}_{0,1}$, $\mathcal{L}_{1,1}$ are integrable [Kac90]. The highest weight vectors $v_{n}\in\mathcal{F}_{\sqrt{2}n}$, $n\in\frac{1}{2}\mathbb{Z}$ in components of decompositions (2.21) are called extremal vectors. It follows from the Theorem 2.15 that $v_{0},v_{1/2}$ are highest weight vectors of $\mathcal{L}_{0,1}$ and $\mathcal{L}_{1,1}$ correspondingly. Other extremal vectors can be found by the formulas
@@
The following formulas for characters follows from Theorem 2.15.
@@ 2.5 Wakimoto module
The $\widehat{\mathfrak{sl}}(2)_{k}$-module $\mathcal{F}_{\lambda,k}$ is called Wakimoto module. See also [FBZ04, Ch.11–12]. The following proposition easily follows from the equality of the characters.
@@ 2.6 Vertex operators
There are (at least) two standard approaches to the definition of vertex operators in conformal field theory. The first one is based on the definition of conformal blocks via coinvariants and operator-state correspondence. In the second approach vertex operators are defined using commutation relation with the algebra generators. We will use both approaches. We start from the first one, mainly following the paper [FM94]. We restrict ourselves to the genus zero conformal block. We fix a global coordinate $t$ on $\mathbb{P}^{1}$.
@@ 2.6.1 Generic representations
This subalgebra is conjugated from $\mathfrak{b}=\langle h,e\rangle$ by the ${\mathrm{e}}^{xf}\in SL(2)$. Let $h_{x},e_{x}$ be generators of $\mathfrak{b}_{x}$ such that $[h_{x},e_{x}]=2e_{x}$, $e_{x}\in[\mathfrak{b}_{x},\mathfrak{b}_{x}]$. For example, for one can take $h_{x}={\mathrm{e}}^{xf}h{\mathrm{e}}^{-xf}=h+2xf$, $e_{x}={\mathrm{e}}^{xf}e{\mathrm{e}}^{-xf}=e-xf-x^{2}f$ for $x\neq\infty$ and $h_{\infty}=-h$, $e_{\infty}=-f$ for $x=\infty$.
@@
$\alpha_{x,z}:\widehat{\mathfrak{sl}}(2)\rightarrow\widehat{\mathfrak{sl}}(2)_{z},$ (2.31)
@@
Of course these formulas do not work if $x$ or $z$ are equal to infinity. So in the neighbourhood of (,)($\infty$,$\infty$)( ∞ , ∞ ) we use another isomorphism
@@
We use the same notation for the corresponding maps between Verma modules $\alpha_{x,z}:\mathcal{M}_{\nu,k}\rightarrow\mathcal{M}_{\nu,k}(x,z)$ mapping $v_{\nu,k}\mapsto v_{\nu,k}(x,z)$.
@@
Let $x_{1},x_{2},x_{3}\in\mathbb{P}^{1}$ and $z_{1},z_{2},z_{3}\in\mathbb{P}^{1}$ be tuples of different points. There is a unique up to constant homomorphism of $\mathfrak{sl}(2)\otimes\mathbb{C}[\mathbb{P}^{1}{\setminus}\{\vec{z}\}]$-modules
@@
We define map $\mathtt{m}_{k}(\vec{x},\vec{z})$ as a composition $\mathtt{m}_{k}(\vec{x},\vec{z})=\mathtt{m}_{k}\circ(\alpha_{x_{1},z_{1}}\otimes\alpha_{x_{2},z_{2}}\otimes\alpha_{x_{3},z_{3}})$
@@
Using $PGL(2)$ action on $\mathbb{P}^{1}$ (global conformal transformations) we can assume that $\vec{z}=\{0,z,\infty\}$, and similarly we can assume that $\vec{x}=\{0,x,\infty\}$ using $SL(2)$ action on the space of Borel subalgebras $\mathbb{P}^{1}=Sl_{2}/B$. We write simply
@@
This Proposition follows from a direct computation (cf. [FM94, Prop. 3.1]). On the other hand this proposition can be considered as a definition of the vertex operator, (see [AY92]).
In order to write down $\mathtt{Y}(v|x,z)$ for arbitrary $v\in\mathcal{M}_{\nu,k}$ we will need notations
$e(x,z)={\mathrm{e}}^{xf_{0}}e(z){\mathrm{e}}^{-xf_{0}}=e(z)-xh(z)-x^{2}f(z),$ (2.40)
$h(x,z)={\mathrm{e}}^{xf_{0}}h(z){\mathrm{e}}^{-xf_{0}}=h(z)+2xf(z),~{}~{}f(x,z)={\mathrm{e}}^{xf_{0}}f(z){\mathrm{e}}^{-xf_{0}}=f(z).$ (2.41)